OpenStudy (anonymous):

Suppose you supply energy to a machine at a rate of 700 w and that the machine converts 560 j into heat every second. At what rate does the machine do work?

OpenStudy (anonymous):

ok eficiency of machine = (Pout/Pin) x 100% eficiency of machine = (500/700) x 100% eficiency of machine = 71.42 %

OpenStudy (anonymous):

efficiency of machine = (Pout/Pin) x 100% efficiency of machine = (500/700) x 100% efficiency of machine = 71.4235 %

OpenStudy (anonymous):

good job @umartouseef :)

OpenStudy (anonymous):

thanx gerryliyana